OpenStudy (barrelracing):

Which of the following is a solution of x2 + 4x + 10?

OpenStudy (barrelracing):

\[A) 2+I \sqrt{6} \]

OpenStudy (barrelracing):

\[B) -2+i \sqrt(24)\]

OpenStudy (barrelracing):

\[c) -2+i \sqrt(6)\]

OpenStudy (barrelracing):

\[d) 2+i \sqrt(24)\]

OpenStudy (barrelracing):

ok all done

OpenStudy (anonymous):

I am working on it at the moment

OpenStudy (barrelracing):

@peachpi

OpenStudy (barrelracing):

@suppa_lova u still there?

OpenStudy (anonymous):

yes I am just a moment

OpenStudy (anonymous):

I would say B is your answer because it is the closest to what I got